The standard mass of an unknown substance is 15.6g. Which of the following set has the least accurate value? Show solutions



Trial A B C


1 15.6 15.8 15.6


2 15.4 14.6 15.5


3 15.6 14.7 15.8


1
Expert's answer
2021-09-28T11:14:12-0400

The least accurate value is the most different from the standard one:


"\delta_1=15.8-15.6=0.2\text g,\\
\delta_2=15.6-14.6=1.0\text g,\\
\delta_3=15.6-14.7=0.9\text g."

Therefore, the second set has the least accurate value.